A book publisher pays $74,200.00 to prepare a book for publication. Total costs are $5.50 per book. If the book is sold to bookstores for $19.50 per copy, how many copies must be sold for the publisher to break even?

Answer by checkley71(8403)

74,200/(19.50-5.50)=74,200/14=5,300 BOOKS MUST BE SOLD TO BREAK EVEN.